OpenStudy (anonymous):

solve the indicated varible -2x+4y=8;x

OpenStudy (anonymous):

-2x = 8 -4y x = 8/-2 -4y/-2 x=-4 +2y

OpenStudy (anonymous):

there we go,lol

OpenStudy (anonymous):

for some reason I kept typing in 9 the first time. sometimes my brain doesn't want to function!

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

x = 2y - 4

OpenStudy (anonymous):

x will be -4